### v1

# Same or Not? Enhancing Visual Perception in Vision-Language Models

Abstract
Vision–language models (VLMs) excel at broad visual understanding but remain coarse-grained, exhibit visual biases, and miss subtle visual details. Existing training corpora reinforce this limitation by emphasizing general recognition (“Is it a cat or a dog?”) over fine-grained perception. To address this, we introduce a new training corpus and task designed to enhance the perceptual abilities of VLMs. Twin is a large-scale dataset of 561,000 image-pair queries that task models to determine whether two visually similar images depict the same object, encouraging attention to nuanced visual cues. The dataset spans a diverse range of everyday objects across contexts, viewpoints, and appearances. Fine-tuning VLMs on Twin yields notable gains in fine-grained recognition, even on unseen domains such as art, animals, plants, and landmarks. To quantify these gains, we introduce Fgvqa, a benchmark suite of 12,000 queries that repurposes fine-grained recognition and retrieval datasets from multiple domains. While existing VLMs struggle on Fgvqa, when fine-tuned on Twin they improve by up to $19.3\%$, without compromising performance on general VQA benchmarks. Finally, our Twin dataset scales favorably with object annotations, and our analysis shows that scale is key to performance. We envision Twin as a drop-in addition to open-source VLM training corpora, advancing perceptual precision of future models. Project webpage: https://glab-caltech.github.io/twin/
(a) The  Twin Dataset
(b) The  Fgvqa Benchmark Suite
Figure 1: (a) Fine-grained visual understanding requires recognizing subtle differences between similar images. We introduce Twin, a large-scale dataset of 561K image-pair queries asking if two images depict the same object, and Fgvqa, a benchmark suite evaluating fine-grained VQA capabilities. (b) Models trained on Twin show improved fine-grained understanding on Fgvqa, even in unseen domains.
1 Introduction
Fine-grained visual understanding – recognizing subtle details within images – is a hallmark of human vision and a desirable goal for machine perception. Modern vision-language models (VLMs) have shown tremendous progress in broad visual reasoning, however their perception remains coarse-grained: they exhibit systematic biases [vlind, visioncheckup, vlmbias, bscore] and overlook subtle details [blink, vlmsblind]. Fig. 1(a) showcases these challenges. While the vacuum cleaners share the same color and brand name, they are distinct instances, distinguishable by the geometry of their dustbins, handle, and base. Qwen2.5-VL [qwen2.5vl], a strong open-source VLM, incorrectly identifies them as identical and reveals flawed reasoning about their visual attributes.
We attribute the limitations of current VLMs in fine-grained perception partly to their training data. Most large-scale image–text corpora emphasize general visual reasoning – such as spatial relations [clevr, marsili2025visual, gqa, cambrian1], common knowledge [mmmu, okvqa], grounding [visualgenome, flickr30k, dci, refcoco, vcr], or mathematical reasoning [mathqa, mathvista, mathverse, math-vision-dataset] – over detailed visual discrimination. While these datasets enable broad understanding, they provide little incentive to attend to subtle, instance-level differences. This imbalance is stark in leading open-data VLMs [perceptionlm, smolvlm], where the majority of their training data does not reward visual distinctions beyond category-level understanding.
We ask: what if VLMs were trained on data that rewards fine-grained visual understanding? To explore this question, we introduce Twin, a large-scale VQA dataset of 561,000561{,}000 instance-centric queries. We name our dataset Twin as it features TWo-image INstance comparisons: each query asks whether two images depict the same physical object instance, requiring attention to subtle visual cues beyond category-level semantics. We source images of household objects across diverse backgrounds, viewpoints, contexts, and lighting conditions, verified by human annotators. We focus on household objects as their images are readily available and they are central to common visual tasks such as robotics and embodied interaction. Fig. 1(a) shows a sample from Twin and illustrates the challenge: distinguishing the vacuum cleaners requires noting differences in form factor and shape, rather than shared brand or color.
To evaluate progress in fine-grained visual perception in VLMs, we introduce Fgvqa, a benchmark suite for fine-grained VQA. While fine-grained recognition tasks have long existed in vision research, they are typically formulated as 1-of-N classification or retrieval problems, making them ill-suited for VLMs. Fgvqa adapts and reformulates existing recognition and retrieval datasets for VLM evaluation, comprising 12,00012{,}000 queries from diverse domains: Met [met] with artworks such as paintings and sculptures, Inquire [inquire] featuring animals and plants from iNaturalist [inaturalist], Cub [cub] with bird species, Ilias [ilias] showcasing retail products, and Landmarks [landmarksv2] containing landmarks such as monuments and architectural structures.
VLMs post-trained on Twin show substantial gains on Fgvqa, as illustrated in Fig. 1(b) which compares the base and post-trained Qwen2.5-VL. Notably, the gains are substantial even on domains that span beyond Twin, e.g. in Inquire and Cub. Through qualitative analyses of model explanations and probing experiments on the vision encoder, we find that VLMs trained on Twin attend more effectively to nuanced, distinguishing visual cues. Finally, our scaling analysis shows that scale is crucial for performance, establishing Twin, which scales in size favorably with the number of object annotations, as a promising path for advancing perceptual precision in VLMs.
2 Related Work
Limitations of VLMs. VLMs trained on large text-image corpora have shown remarkable capabilities in image understanding. Here, proprietary systems [gpt-5, gpt4, gemini2.5, claude] lead in performance, while open-source alternatives [vila, deepseekvl2, llama3, molmo, moondream2, smolvlm, gemma] are narrowing the gap. Despite these advances, VLMs exhibit systematic biases [vlmbias, vlind, bscore, visioncheckup], struggle with fine-grained visual perception [blink, vlmsblind], and often hallucinate details [pope, eyeswideshut, autohallusion, beaf]. To address these challenges, we introduce a large-scale VQA dataset to enhance fine-grained visual understanding in VLMs along with a benchmark suite to evaluate this capability.
VLM Training Data. While most proprietary VLMs rely on closed-source training data [gpt-5, gpt4, claude, gemini2.5, gemini], recent efforts have prioritized open-data alternatives. Molmo [molmo] introduced an open-data model competitive with proprietary systems, while SmolVLM [smolvlm] and PerceptionLM [perceptionlm] scale this paradigm using diverse datasets spanning captioning [coco, nocaps, young2014image, docci, altogether], grounding [visualgenome, flickr30k, dci, refcoco, vcr], spatial reasoning [blink, clevr, vsr], math [intergps, gsm8k, mathqa], and chart/diagram understanding [chartqa, charttotext, ai2d, vistext, charxiv]. Most relevant to us are SpotTheDiff [spotthediff] and Birds-to-Words [birdstowords], which pair images with text describing differences. In contrast, Twin is more comprehensive and larger in size: Twin features both matching and non-matching pairs, enabling reasoning about sameness and difference, distinctions in Twin are finer and pertain to subtle cues, Twin is far larger (561561K pairs vs. 1313K of SpotTheDiff and 3.33.3K of Birds-to-Words), and it spans diverse object types beyond a single domain such as birds. A detailed comparison is in the Appendix.
Fine-grained Recognition categorizes images into specific subgroups [finegrainedsurvey], with applications in classifying plants [inaturalist, oxfordflowers, inquire], birds [cub2011, nabirds], vehicles [fgvcaircrafts, stanfordcars] and recently, retail products [imaterialist, products10k, aliproducts, ilias]. These tasks are typically framed as 1-of-N classification or retrieval, making them ill-suited for evaluating VLMs. We repurpose popular benchmarks [ilias, met, inquire, cub, landmarksv2] and design a task to evaluate the fine-grained VQA abilities of VLMs.
VLM Post-training. Post-training refines VLMs by instilling specialized capabilities and aligning them more closely with human intent. Supervised fine-tuning (SFT) [star, s1, vocot] and distillation [llava-cot, visprogdistill, ical] transfer task-specific skills from annotated data or teacher models, but often cause forgetting and reduced generalization [rlsrazor, sftmemorizesrlgeneralizes]. Reinforcement learning mitigates these issues, advancing instruction-following [rlhf, dpo, qwen3], reasoning [deepseek-r1, deepseekvl2, drgrpo, qwen3, gemini2.5], and spatial understanding [vigorl, grit, marsili2025no]. Beyond post-training methodology, the choice of data plays a critical role in driving post-training success. We demonstrate that reinforcement learning on our fine-grained Twin dataset increases the perceptual precision of VLMs while preserving their performance on core image–language tasks.
Figure 2: (a) Twin is a large-scale VQA dataset for fine-grained visual understanding, where VLMs determine whether two images depict the same instance. Twin contains 561561K pairwise VQA queries across $1,836$ object instances, spanning $36$ categories of common objects and over 2222K images. (b) Scalability of the pairwise construction: from $1,836$ instances, Twin yields 561561K pairwise queries.
3 The Twin Dataset
We present Twin, a large-scale VQA dataset for advancing fine-grained visual understanding in VLMs. Existing training corpora emphasize broad visual understanding – spatial relations, common knowledge, and category-level recognition – offering little supervision for fine perception. As shown in Fig. 1(a), current VLMs often miss subtle cues essential for reliable perception. To address this, Twin introduces 561,000561{,}000 instance-centric queries where models are tasked to judge whether two similar looking images depict the same object instance. This design rewards attention to nuanced, instance-level details such as shape, texture, and part geometry, going beyond category-level understanding.  Fig. 2 shows samples from Twin and its statistics.
3.1 Dataset Construction
Task Definition. In our task, fine-grained visual understanding requires distinguishing object instances – recognizing when two images show the same object rather than the same category. We define an instance as a set of images of the same physical object under varied viewpoints, lighting, and backgrounds, consistent with prior work [ilias, met, scalablerecognitionvocabtree, landmarksv2, image-retrieval-deep-local]. From this definition, we design a VQA task where a VLM receives two images and determines if they depict the same instance. Unlike category-level recognition (e.g., cat vs dog), our task demands sensitivity to fine, instance-specific cues – shape, texture, and part geometry – that distinguish similar objects. Fig. 1(a) shows the setup; VQA prompts are in the Appendix.
Figure 3: Fgvqa is a suite of fine-grained VQA benchmarks spanning retail products (Twin, Ilias [ilias]); animals and plants (Inquire [inquire]); Landmarks [landmarksv2]; birds (Cub [cub]); and art (Met [met]). We include two query types: pair (top row), where a VLM judges if two images depict the same instance, species, or landmark, and multi where it counts how many images match a reference.
Sourcing Instances. We source object instances across diverse categories from Amazon Reviews [amazon-reviews]. For each object instance, we collect images with varied contexts, viewpoints, and lighting to ensure visual diversity. Human annotators verify that each image clearly shows and identifies the object. We generate positive pairs by sampling from images of the same instance. Fig. 2(a) shows examples of positive pairs along with the dataset’s distribution.
Hard Negative Pairs. A balanced dataset for fine-grained understanding requires both positive and negative pairs. If all examples were positive, the task would be trivial. Likewise, random negatives are often too easy (e.g., a mug paired with a fan). We therefore focus on hard negatives – distinct objects that appear similar. We collect these hard negatives with the help of human annotators. First, using CLIP [clip], we compute pairwise cosine similarities between image embeddings to shortlist visually similar candidates, from which human annotators select the final pairs. Examples of such hard negative pairs are shown in Fig. 2(a), including two white vases with slightly different shapes and two white chairs differing in texture and armrest design.
Generated Negatives. Collecting hard negatives is costly and difficult to scale, so we augment Twin with synthetic negatives using personalized image generation, DreamBooth [dreambooth]. Given images of an instance, DreamBooth produces similar variants that retain overall appearance but alter fine details, simulating hard negatives without extra data collection. The rightmost speaker in Fig. 2(a) demonstrates this: its color and geometry match the left one, but it omits the red logo and lighter contours. Each synthetic image is verified by human annotators. Details on image generation and additional examples are in the Appendix. We validate the impact of these generated negatives in our experiments.
Dataset Statistics. Twin features 561561K pairwise VQA queries, with 123123K positive and 438438K negative samples, spanning 22,15722{,}157 unique images of 1,8361{,}836 object instances, with 5,2885{,}288 generated images. We plot the number of pairwise queries vs the number of object instances in Fig. 2(b), highlighting the scalability of our pairwise formulation: Twin’s size scales favorably with the number of object instances. Our scaling analysis in Sec. 5.2 demonstrates that collecting Twin at scale is crucial for performance.
3.2 Training VLMs on Twin
Figure 4: Training VLMs on Twin. We train VLMs using RL on Twin. Reward is computed by comparing the predicted answer with the ground truth pair assignment.
To improve fine-grained understanding in VLMs and evaluate the impact of our new dataset, we post-train existing models on Twin. Our task requires recognizing subtle attributes to distinguish similar instances, and we hypothesize that optimizing for this task enhances broader fine-grained abilities. We post-train with RL as it is shown to improve model capabilities while preserving prior skills [sftmemorizesrlgeneralizes, rlsrazor, huan2025does, unveilingcompositiongap]. Fig. 4 outlines our model training approach.
Setup. Given image pairs $(I_{1},I_{2})$ with ground truth label $y\in\{\text{yes, no}\}$, a VLM $\pi_{\theta}$, parametrized by $\theta$, is prompted to produce a textual explanation and a final answer $\hat{y}$ whether both images depict the same instance. We use a binary outcome reward comparing prediction and ground truth: $R(y,\hat{y})=\mathbf{1}_{\{y=\hat{y}\}}$. Importantly, supervision relies only on pairwise assignments, without any descriptive textual annotations.
Optimization. We tune our VLM $\pi_{\theta}$ using GRPO [deepseekmath] as it balances maximizing expected advantages while preventing drift from the pre-trained VLM. Details are in the Appendix.
Implementation. We train Qwen2.5-VL-3B-Instruct [qwen2.5vl] and InternVL3.5-1B-Instruct [internvl3] on Twin, chosen as leading open-source VLMs at the 3B and 1B scales. Training is end-to-end on $4$ A100 GPUs for $1$ epoch with a batch size of $480$, group size $5$, and learning rate $10^{-6}$. We use verl [verl] for optimization and balance positive and negative queries in each batch. Hyperparameters are in the Appendix.
4 The Fgvqa Benchmark Suite
Despite their strong general visual abilities, current VLMs struggle with fine-grained understanding [vlmbias, vlmsblind]. Existing VQA benchmarks emphasize broad image understanding, leaving this capability largely unevaluated. To this end, we introduce Fgvqa, a suite of fine-grained VQA benchmarks that measure how well VLMs identify subtle visual cues. We show examples in Fig. 3 and provide more details below.
Benchmarks. Fine-grained understanding is a general skill – models attuned to subtle differences should generalize across domains. To evaluate this, we repurpose recognition and retrieval datasets, totaling 12,00012{,}000 queries spanning diverse domains: Met [met] (artworks such as paintings and sculptures), Ilias [ilias] (retail products), Landmarks [landmarksv2] (monuments), Cub [cub] (bird species), Inquire [inquire] (animals and plants from iNaturalist [inaturalist]), and Twin-Eval, collected like Twin but with unseen instances. The breadth of Fgvqa enables assessment of cross-domain generalization.
Twin-Eval Ilias [ilias] Landmarks [landmarksv2] Met [met] Cub [cub] Inquire [inquire]
Mean Pair Multi Total Pair Multi Total Pair Multi Total Pair Multi Total Pair Multi Total Pair Multi Total
[…]
GPT4o [gpt4] $86.1$ $92.2$ $71.2$ $81.7$ $95.0$ $95.5$ $95.3$ $87.3$ $74.7$ $81.0$ $86.3$ $72.9$ $79.6$ $91.7$ $86.0$ $88.9$ $94.3$ $86.3$ $90.3$
Gemini 2.5 Flash [gemini2.5] $82.2$ $89.1$ $68.1$ $78.6$ $92.2$ $90.5$ $91.4$ $85.4$ $77.0$ $81.2$ $82.6$ $71.9$ $77.3$ $84.1$ $77.9$ $81.0$ $88.3$ $79.4$ $83.9$
[…]
PerceptionLM 3B [perceptionlm] $37.8$ $46.9$ $25.2$ $36.1$ $49.6$ $25.0$ $37.3$ $51.5$ $25.7$ $38.6$ $50.4$ $26.1$ $38.3$ $49.6$ $24.9$ $37.3$ $52.1$ $26.1$ $39.1$
Moondream 2 2B [moondream2] $45.9$ $64.5$ $30.0$ $47.3$ $58.7$ $30.3$ $44.5$ $59.8$ $25.2$ $42.5$ $84.3$ $27.0$ $55.7$ $50.6$ $43.3$ $47.0$ $51.9$ $25.1$ $38.5$
Gemma3 4B [gemma3] $68.7$ $68.1$ $39.5$ $53.8$ $88.4$ $46.5$ $67.5$ $91.2$ $55.2$ $73.2$ $87.9$ $57.5$ $72.7$ $91.2$ $55.2$ $73.2$ $90.0$ $48.7$ $69.4$
SmolVLM2 2.2B [smolvlm] $46.9$ $54.7$ $25.1$ $39.9$ $52.2$ $29.7$ $41.0$ $54.1$ $29.6$ $41.9$ $73.9$ $27.9$ $50.9$ $81.5$ $35.0$ $58.3$ $73.0$ $25.7$ $49.4$
[…]
Qwen2.5-VL 3B [qwen2.5vl] $53.0$ $70.4$ $29.8$ $50.1$ $54.5$ $32.5$ $43.5$ $79.5$ $28.2$ $53.9$ $81.6$ $29.4$ $55.5$ $84.0$ $37.4$ $60.7$ $74.4$ $34.3$ $54.4$
+ Twin 67.0+14.0 88.9+18.5 45.7+15.9 67.3+17.2 79.9+25.4 43.7+11.2 61.8+18.3 80.8+1.3 34.9+6.7 57.9+4.0 85.4+3.8 46.5+17.1 66.0+10.5 92.5+8.5 57.7+20.3 75.1+14.4 92.7+18.3 54.6+20.3 73.7+19.3
InternVL3.5 1B [internvl3] $51.3$ $67.2$ $25.9$ $46.6$ $60.0$ $39.1$ $49.6$ $54.1$ $33.5$ $43.8$ $73.9$ $47.1$ $60.5$ $56.1$ $40.5$ $48.3$ $80.4$ $37.9$ $59.2$
[…]
Table 1: Accuracy (%) on Fgvqa. For Pair queries, VLMs are shown two images and asked if they show the same instance. On multi queries, VLMs are shown a reference image and three candidates and asked how many show the same instance as the reference. We report accuracy on each constituent dataset separately and the mean. Training on Twin improves performance, even on domains not in Twin.
Query types. For Fgvqa, we construct two query types per dataset. Pair queries show two images and ask whether they depict the same instance, artwork, or species. Multi queries provide a reference image and three candidates, and ask how many match the reference. Each dataset includes $1,000$ balanced examples per type: pair queries split evenly among positive and negative cases, while multi queries distribute uniformly across answer counts ($250$ each from 0-33). Examples are shown in Fig. 3, Fig. 5, and the Appendix.
Applicability. Fgvqa provides a general framework for evaluating fine-grained recognition in VLMs. First, it probes visual perception – testing whether models can distinguish subtle differences that support accurate reasoning. Second, the pairwise yes/no format aligns with recognition and retrieval tasks, since both 1-of-N classification and retrieval can be cast as a sequence of pairwise comparisons. Together, these properties make Fgvqa a practical, domain-general suite for analyzing perception in VLMs.
5 Experiments
We conduct extensive experiments, both quantitative and qualitative, to evaluate the utility of Twin as a training dataset and task for improving fine-grained understanding. Our goals are threefold: (1) establish baseline performance on Fgvqa, our new fine-grained VQA benchmark suite that repurposes existing datasets, (2) assess if training on Twin improves perception across domains, and (3) ablate the decisions made in our dataset construction and model training.
Figure 5: Outputs on Fgvqa for base VLMs and variants trained on Twin. For each example, we show the source dataset, images, and both model predictions. We include both pair queries (a-d) and multi queries (e-h). Incorrect reasoning is highlighted in red and correct reasoning in green. The red circle highlights a hard-to-see object. Training on Twin enhances fine-grained grounding, leading to increased attention to part geometry and texture over the base model, even for domains that differ from the household objects in Twin.
Baselines. We compare to the following models:
(1) Proprietary VLMs: We establish baseline performance on Fgvqa for GPT4o [gpt4] and Gemini 2.5 Flash [gemini2.5].
(2) Open-Source VLMs: We report leading open-source models PerceptionLM-3B [perceptionlm], Moondream2-2B [moondream2], Gemma3-4B [gemma3], and SmolVLM2-2.2B [smolvlm].
(3) Direct Comparisons: We fine-tune Qwen2.5-VL 3B Instruct [qwen2.5vl] and InternVL3.5 1B Instruct [internvl3] to evaluate the value of Twin across architectures and scales. Each base model is compared to its trained variant, denoted “+ Twin”.
5.1 Main Results
Sec. 4 shows results on Fgvqa. Model variants trained on Twin are evaluated zero-shot on all datasets except Twin-Eval. All other models are evaluated zero-shot. All models are prompted without in-context examples. We summarize our findings below.
Training on Twin improves fine-grained understanding. Our direct comparisons show substantial gains from training on Twin. Qwen2.5-VL (3B) achieves strong improvements across datasets: $18.3\%$ on Ilias ($43.5\%$ to $61.8\%$) and $17.2\%$ on Twin-Eval ($50.1\%$ to $67.3\%$). InternVL3.5 (1B), despite being smaller and thus less expressive, still shows consistent gains from post-training on Twin: $3.8\%$ on Cub ($48.3\%$ to $52.1\%$) and $3.0\%$ on Landmarks ($43.8\%$ to $46.8\%$). These consistent improvements across both models indicate that Twin benefits VLMs of different scales and architectures. Importantly, we observe improvement on multi queries, which differ in structure from our training data, suggesting improved fine-grained perception rather than task overfitting.
Improvements transfer to unseen domains. Training on Twin, which features everyday household objects, improves performance even in domains absent from its training distribution. Substantial gains are observed on animal and plant species (Inquire & Cub), and for art and landmarks (Met & Landmarks), which are distinct from objects in Twin. With Qwen2.5-VL, Inquire accuracy rises by $19.3\%$ ($54.4\%$ to $73.7\%$), Cub by $14.4\%$ ($60.7\%$ to $75.1\%$), and Landmarks by $4\%$ ($53.9\%$ to $57.9\%$).
Qualitative Examples. We present qualitative results in Fig. 5, comparing baseline models with their counterparts post-trained on Twin. The textual explanations reveal which visual cues models rely on, showing how Twin enhances fine-grained reasoning. In (a), the trained Qwen2.5-VL distinguishes subtle color differences in the numbers and clock hands. In (c), it correctly identifies shared classical columns. In (e), it matches the reference card under challenging lighting in Image 2. In (g), the base model incorrectly flags Image 2 as a different knife, while the trained variant recognizes consistent blade texture and handle material. In (h), only the trained model notices color and design differences in the synthetic negative speaker (Image 3).
These improvements extend to out-of-distribution examples. In (b), the post-trained model identifies the bird by attending to its black head, white body, and orange beak. In (d), it notes differences in date and color scheme. In (f), it detects color differences between the reference bird and Image 3. These improved reasoning descriptions confirm that Twin enhances understanding beyond the training domain.
Comparison to open-source VLMs. Among open-source baselines, Gemma3 performs best across benchmarks: $53.8\%$ on Twin-Eval, $69.4\%$ on Inquire, and $72.7\%$ on Ilias. However, as Gemma3 does not support training with Flash Attention [flashattention2, gemma-flash-attn1, gemma-flash-attn2], fine-tuning it in PyTorch would be prohibitively expensive. Therefore, we train the second-best open-source VLM, Qwen2.5-VL, on Twin. Our trained Qwen2.5-VL model (+ Twin) outperforms Gemma3 on Twin-Eval ($67.3\%$ vs $53.8\%$), Inquire ($73.7\%$ vs $69.4\%$), and Cub ($75.1\%$ vs $73.2\%$). Gemma3 outperforms our trained Qwen2.5-VL model on Met ($72.7\%$ vs $66.0\%$) and Landmarks ($73.2\%$ vs $57.9\%$) – the datasets where we see the least improvement from training.
Comparison to proprietary VLMs. We also benchmark proprietary models on Fgvqa. GPT4o and Gemini-2.5-Flash outperform all open-source baselines ($81.7\%$ for GPT4o vs $53.8\%$ for Gemma3 on Twin-Eval; $91.4\%$ for Gemini 2.5 Flash vs $67.5\%$ for Gemma3 on Ilias). These models are likely far larger and trained on proprietary data. Training on Twin narrows this gap: Qwen2.5-VL + Twin reaches $73.7\%$ on Inquire, reducing the gap from $29.5\%$ to $10.2\%$, showing that targeted fine-grained training substantially closes the gap with proprietary models.
5.2 Ablations and Analysis
We now analyze factors driving improved performance on Fgvqa. First, we study the Twin dataset and examining choices made during its construction, namely the challenging negative pairs (real and synthetic) in Tab. 2 and data scale in Fig. 6. Next, we compare post-training paradigms on Twin, supervised fine-tuning vs reinforcement learning, in Tab. 3. Finally, we probe the vision encoder to showcase that training on Twin improves underlying representations in Tab. 5. We summarize our findings below.
Mean Twin-Eval Ilias Landmarks Met Cub Inquire
Qwen2.5-VL 3B 53.0 50.1 43.5 53.9 55.5 60.7 54.4
+ Twin w/o Hard Neg. 58.6 51.1 53.1 57.9 60.0 68.5 60.9
+ Twin 63.1 65.3 58.4 54.9 64.4 66.9 68.7
Table 2: Impact of hard negatives. Models in rows 2 and 3 are trained on the Twin variant of their corresponding row. The hard negatives in Twin significantly improve performance on Fgvqa.
Hard Negatives. We evaluate the contribution of hard negatives to Fgvqa performance in Tab. 2. We post-train Qwen2.5-VL 3B on two variants: (1) Twin, a 250250K-sample subset matching the original ratios of positive/negative and real/generated images—our final setting; and (2) Twin w/o Hard Neg., where all hard negatives are replaced with random negatives. This controlled comparison isolates the value of including hard negatives in Twin. We report total accuracy ($\%$) on each Fgvqa dataset separately, as well as the mean across datasets.
Tab. 2 shows that even without hard negatives, our task yields modest improvements over the base model ($51.1\%$ vs $50.1\%$ on Twin-Eval; $60.0\%$ vs $55.5\%$ on Met). However, incorporating hard negatives improves performance by $4.5\%$ on average ($63.1\%$ vs $58.6\%$). The gains are most pronounced on in-domain datasets like Twin-Eval ($65.3\%$ vs $51.1\%$) and Ilias ($58.4\%$ vs $53.1\%$), but remain substantial on out-of-distribution datasets like Inquire ($68.7\%$ vs $60.9\%$). These results validate our decision to collect human-verified hard negatives, demonstrating they are critical for developing fine-grained visual understanding.
Scaling Analysis. Fig. 6 shows the impact of data scale for Twin. We train a Qwen2.5-VL-Instruct 3B model with $M=\{5$K,50,50K,250,250K,561,561K}\} image pairs and plot accuracy on Fgvqa. We sample $M$ with the same proportions of positive/negative pairs and real/generated images as our full dataset. Performance improves consistently across all datasets from 55K to 561561K samples (e.g. $48.5\%$ to $67.3\%$ on Twin and $44.2\%$ to $61.8\%$ on Ilias), reinforcing our decision to collect Twin at scale. Notably, scale also improves performance on domains not represented in Twin: Inquire improves from $53.3\%$ to $73.7\%$ and Cub from $62.2\%$ to $75.1\%$ as training data scales from 55K to 561561K.
Figure 6: Effect of training set size on Fgvqa accuracy (%).
Mean Twin-Eval Ilias Landmarks Met Cub Inquire
RL 57.5 64.0 50.3 48.5 58.7 65.0 58.7
SFT 53.8 61.2 53.6 47.6 52.3 53.9 54.2
Table 3: SFT vs. RL. We tune Qwen2.5-VL 3B on a subset of Twin with SFT and RL. We report total accuracy (%) on Fgvqa.
SFT vs. RL. We compare post-training with reinforcement learning (RL) versus supervised fine-tuning (SFT) in Tab. 3. Our model produces both a chain-of-thought (CoT) explanation and a final yes/no prediction. RL optimizes a reward based solely on the final answer correctness, while SFT performs supervision on all output tokens, including both CoT and final prediction. To obtain high-quality supervision for SFT, we collect responses from Gemini-2.5-Flash [gemini2.5] on a subset of Twin and retain only those where Gemini is correct. We then post-train Qwen2.5-VL on these 136136K pairs with both SFT and RL. Details for SFT are in the Appendix. Although both methods achieve similar performance on Twin-Eval ($61.2\%$ with SFT vs $64.0\%$ with RL), the RL variant outperforms on out-of-distribution datasets ($53.9\%$ with SFT vs $65.0\%$ with RL on Cub; $52.3\%$ with SFT vs. $58.7\%$ with RL on Met). These results corroborate prior findings that RL better preserves generalization during post-training than SFT [sftmemorizesrlgeneralizes, rlsrazor, huan2025does, unveilingcompositiongap].
Realworld [realworldqa] Pope [pope] Cv-Bench [cambrian1] Nlvr2 [nlvr2] Ai2d [ai2d] Seed [seedbench] ChartQA [chartqa] TextVQA [textvqa] Blink [blink] Mmmu [mmmu] VlmBias [vlmbias]
[…]
Qwen2.5 VL 3B [qwen2.5vl] 59.8 87.6 66.0 79.8 78.6 75.0 83.4 79.4 47.3 47.1 17.2
+ Twin 61.8 88.2 66.6 81.2 79.4 75.1 84.1 77.9 47.5 46.9 17.4
InternVL3.5 1B [internvl3] 52.2 86.3 59.8 73.6 67.1 70.9 63.0 55.9 39.3 40.4 17.4
+ Twin 52.6 86.3 60.2 73.8 67.2 70.8 63.2 55.7 39.7 40.1 17.4
Table 4: Accuracy (%) on general VQA benchmarks. VLMs post-trained on Twin retain performance on general VQA benchmarks.
Pets [oxfordpets] Sun397 [sun397] Cub [cub] Cifar100 [cifar100]
Lin. KNN Lin. KNN Lin. KNN Lin. KNN
Qwen2.5 VL 3B [qwen2.5vl] 75.0 72.4 64.4 70.4 72.3 77.4 68.5 71.8
+ Twin 79.1 74.5 66.4 71.8 72.4 77.7 69.6 73.2
Table 5: Encoder probing results. We evaluate Qwen2.5-VL’s vision encoder via linear probing and KNN. Post-training on Twin yields embeddings better suited for fine-grained classification.
Encoder Probing. We probe the underlying vision encoder of Qwen2.5-VL to examine whether training on Twin improves its visual representations. Unlike contrastive models (e.g. CLIP [clip]), this encoder was not trained to produce embeddings useful in isolation. Nevertheless, we apply traditional probing methods to assess representation quality and determine if post-training on Twin improves representation ability of vision encoders in VLMs. In Tab. 5, we report accuracy on fine-grained classification datasets covering various domains: Pets [oxfordpets] with cat and dog species, Sun397 [sun397] with real world scenes, Cub [cub] with bird species, and Cifar100 [cifar100] with everyday objects and animals. We follow [perceptionencoder] and evaluate two settings: a linear probe (Lin.) trained on embeddings from each dataset’s training split and then frozen for validation, and K-nearest neighbor retrieval (KNN), where test queries are matched against embeddings from the corresponding training split.
From Tab. 5, we find that post-training on Twin yields consistent improvements across fine-grained classification datasets: $4.1\%$ with a linear probe on Pets ($75.0\%$ to $79.1\%$), $1.4\%$ with KNN on Cifar100 ($71.8\%$ to $73.2\%$), and $2\%$ with a linear probe on Sun397 ($64.4\%$ to $66.4\%$). The improvement to probing accuracy across datasets suggest that by forcing VLMs to detect subtle differences in objects, post-training on Twin yields embeddings better suited for fine-grianed understanding.
5.3 Additional Results
To ensure post-training on Twin does not erode general visual reasoning, we report results on popular VQA benchmarks in Sec. 5.2. These benchmarks span common sense understanding (Seed [seedbench], Mmmu [mmmu], Nlvr2 [nlvr2]), spatial reasoning and grounding (Realworld [realworldqa], Cv-Bench [cambrian1], Blink [blink]), hallucination and bias (Pope [pope], VlmBias [vlmbias]), OCR (TextVQA [textvqa]), and chart/diagram understanding (ChartQA [chartqa], Ai2d [ai2d]). Although not our main focus, models post-trained on Twin retain performance on general VQA benchmarks, with small improvements on most for both Qwen2.5-VL ($1.4\%$ on Nlvr2; $0.8\%$ on Ai2d) and InternVL3.5 ($0.4\%$ on Realworld; $0.4\%$ on Cv-Bench).
Mmlu [mmlu] HellaSwag [hellaswag] Gsm8K [gsm8k]
Qwen2.5 VL 3B [qwen2.5vl] 64.7 66.4 70.8
+ Twin 64.6 66.4 68.8
InternVL3.5 1B [internvl3] 49.0 48.8 57.8
+ Twin 49.0 48.9 58.1
Table 6: Accuracy (%) on text-QA benchmarks. Post-training on Twin with RL leaves text-QA performance largely unaffected.
Beyond vision, Tab. 6 reports performance on popular text-only benchmarks covering different skills: Mmlu [mmlu] for common knowledge, HellaSwag [hellaswag] for sentence completion, and Gsm8K [gsm8k] for mathematical reasoning. We find that post-training leaves performance largely unchanged on these benchmarks: $64.7\%$ vs $64.6\%$ for Qwen2.5-VL on Mmlu; $57.8\%$ vs $58.1\%$ for InternVL3.5 on Gsm8K. These results reaffirm prior findings that RL-based tuning preserves model skills [rlsrazor, sftmemorizesrlgeneralizes].
6 Limitations & Future Work
We introduce Twin, a large-scale VQA dataset of 561,000561{,}000 queries designed for improving fine-grained perception in VLMs. Each query in Twin asks whether two images depict the same object instance, forcing VLMs to perceive subtle visual cues to answer correctly. Twin includes human-verified hard negative pairs, which we determine are crucial in our experiments. To measure progress on fine-grained understanding, we additionally introduce Fgvqa, a benchmark suite for precise visual understanding across a wide range of domains. Current open-source VLMs struggle on Fgvqa, but post-training them on Twin substantially improves fine-grained reasoning, even on unseen domains. Our scaling analysis shows that collecting Twin at scale is key for improved performance. We include a thorough failure analysis of post-trained models in the Appendix. There is an extensive list of future work to consider:
  • We explore GRPO post-training with rewards applied only on final answers. More structured rewards, e.g. using multi-modal verifiers, may provide a richer learning signal and lead to improved perception capabilities.
  • Our experiments show hard negatives in Twin are essential for effective learning. Exploring automated techniques to mine even harder negatives, e.g. with the help of model-in-the-loop data engines, is a promising direction.
  • Twin includes substantial viewpoint variation, requiring models to reason over fine-grained part geometries that may appear different across image pairs. Integrating 3D representations as an explicit means of encoding spatial structure may further improve performance on the task.
We envision Twin as a drop-in addition to VLM training corpora and hope Fgvqa serves as a benchmark suite for measuring progress in fine-grained understanding. To support future research in this direction, we will release all data, code, and post-trained models.
Acknowledgments
We thank Aadarsh Sahoo, Ilona Demler, and Ziqi Ma for their feedback on the project. The project is funded by Meta through the LLM evaluation research grant and partly through Caltech’s CAST program. We also thank Google’s Gemma Academic program for granting us API credits for their LLMs.
Supplementary Material
Figure 7: Additional examples from Fgvqa. We show examples of images in pair and multi queries. Fgvqa spans six datasets across a wide range of domains, creating a challenging suite for assessing cross-domain fine-grained VQA in VLMs.
Appendix A Additional details on Fgvqa
We provide additional details on Fgvqa, our new benchmark suite for evaluating fine-grained VQA. We describe the datasets included in Fgvqa in Sec. A.1. We illustrate additional examples in Fig. 7, and describe all prompts used in Sec. A.2. We report exact match accuracy for all benchmarks on both pair and multi queries.
A.1 Evaluation Benchmarks
We describe the provenance of each benchmark in Fgvqa below.
Twin-Eval is the evaluation set of Twin. It is collected identically to Twin (Sec. 3), but features distinct instances and images from Twin.
ILIAS [ilias] is a large-scale test dataset of instance-level image retrieval. It predominantly features images of retail products taken in various contexts, backgrounds, and lighting. For both pair and multi queries, we source images from the core_db split, using the __key__ field to determine instance identity.
Google Landmarks v2 (Landmarks) [landmarksv2] is a landmark recognition dataset featuring human-made and natural landmarks. The dataset has been used in both classification and retrieval settings. To ensure sufficient images per landmark to for multi queries, we source images from the train split and use the label field to determine landmark identity.
MET [met] is an image retrieval dataset featuring artwork from the Metropolitan Museum of Art in New York. The dataset features images of the same art piece or sculpture from varying viewpoints, emphasizing multi-view consistency in retrieval. We use the mini_met set to construct both pair and multi queries.
CUB [cub] is a fine-grained classification dataset that focuses on identifying bird species from images. We use the test split for all queries and create pair and multi queries based on the specie of bird.
Inquire [inquire] is a benchmark for natural world image retrieval, featuring images of animal and plant species sourced from iNaturalist [inaturalist]. We source images from the validation split and use the inat24_species_name field to determine instance identity for both pair and multi queries.
A.2 Fgvqa Prompts
We include all prompts used on all datasets for both pair and multi queries in Figs. 12 to 17. All prompts are identical in format, except for the domain-specific instance names (e.g. bird/object/artwork/species). Additionally, for benchmarks featuring object instances (Twin-Eval, Ilias [ilias]), we found it necessary to provide the definition of an object instance as models were incorrectly labeling different objects of the same-category (e.g. two earbuds of different colors) as the same.
Mean Twin-Eval Ilias Landmarks Met Cub Inquire
Human 92.9 90.0 100.0 82.5 93.8 92.5 98.8
Table 7: Human Evaluations on Fgvqa.
A.3 Fgvqa Human Evaluation
We establish human performance on Fgvqa in Tab. 7. We task four human annotators with repsonding to a subset of 20 queries per benchmark in Fgvqa. We report total accuracy (%) on each dataset, averaged across evaluator. We find that human evaluators significantly outperform open-source VLMs, suggesting ample future work is needed in fine-grained VQA.
Appendix B Additional details on Twin
We provide further details on the collection of Twin, including instance sourcing, de-duplication, hard negative pair assignment, and details on generating synthetic negatives.
B.1 Instance Sourcing
We used the Amazon Reviews dataset [amazon-reviews] to extract a large set of product listings, which would serve as candidate objects, and associated noisy review images. We categorized these objects into supercategories based on their associated category in the Amazon store (e.g. “Electronics”) and further into categories based on the product title (e.g. “Speaker”). We aimed to maximize diversity in our instances and thus selected 36 categories to draw from. We subsequently tasked human annotators with removing products that either lacked visual consistency (e.g. the same product being sold in multiple colors) or did not adhere to the object category (e.g. “earbud replacement tips” instead of “earbuds”).
Figure 8: Additional Synthetic Negatives. For each pair, we show a real image of the object instance and a synthetic image of the same instance generated with Dreambooth [dreambooth]. We find the synthetic negatives capture the “gist” of the instance, but are not faithful to fine details such as color, geometry, and texture – making them a strong additional set of hard negative pairs.
B.2 De-Duplication
To ensure the correctness of the pairwise labels in Twin, we must ensure that object instances are unique. If object instances are not unique, then two of the same objects could incorrectly be listed as negative pairs. However, on Amazon duplicate listings are common. We therefore merged duplicate objects into the same instance. To this end, we used a pre-trained CLIP model to compare the images of each product with the images of all other products via cosine similarity. We flagged the $k$ images with the highest similarity as potential duplicates, setting $k$ equal to the number of images for the product. We then tasked human annotators with manually inspecting the flagged images. If a duplicate was found, we merged the two products. The resulting object instances after de-duplication are used to produce the 561561K instances of Twin.
B.3 Hard Negative Pair Assignment
Once we have collected a set of object instances, we construct hard negative pairs – pairs of distinct objects that appear similar. We collect these hard negatives with the help of human annotators. First, using CLIP [clip], for each object instance, we compute pairwise cosine similarities between the image embeddings of that instance and all other instances. We use these similarities to shortlist $k$ visually similar candidates for each instance, with $k$ being a random number between $1-2\times$ the number of positive images of that instance. Human annotators then select the final pairs, omitting pairs they deem to be too easy.
B.4 Additional Details on Generated Negatives
We augment the set of negative pairs with additional synthetically generated images. Augmenting with generated negatives allows us to produce additional negative pairs without extra data collection. We follow the procedure outlined in Dreambooth [dreambooth]. For each object instance, we sample 3 to 5 images for each instance to serve as grounding images. We set the initializer tokens to the generic category associated with the instance (e.g. “vase”, “headphones”). Our prompts to Dreambooth are: [“an image of {category}”, “similar to {category}”, “a picture of{category}, “show me a {category}, “here is a {category}”] replacing {category} with the category of the instance. Finally, we task human annotators with validating the generated samples. Annotators are asked to remove generations that either do not appear similar to the instance or are indistinguishable from the instance – yielding a set of hard negative pairs.
We show additional examples of synthetic negative pairs in Fig. 8. For each pair of images, we show the real image on the left and the paired synthetic image on the right. We find that synthetic negatives faithfully capture the “gist” and overall appearance of the object instance, but they fail to capture nuanced details such as part color and geometry, texture, and brand text/logos. These subtle differences in appearances yield an augmenting set of hard negative pairs.
Appendix C Additional Training Details
We include additional details for all model training. In Sec. C.1, we provide an overview of GRPO [deepseekmath], our main post-training method, and our training setup. We provide additional details on supervised fine-tuning (Sec. 5.2) in Sec. C.2. We include all hyperparameters in Tab. 8 and Tab. 9.
C.1 GRPO
To post-train VLMs on Twin, we use Group Relative Policy Optimization (GRPO; [deepseekmath]). We provide an overview of this method below.
Setup. Given a pair of images $(I_{1},I_{2})$ with ground-truth pairwise label $y\in\{\text{yes},\text{no}\}$, our base VLM $\pi_{\theta}$, parametrized by $\theta$, is prompted to produce a reasoning explanation $r$ and final answer $\hat{y}$ whether both images depict the same instance. We wish to find $\pi_{\theta^{*}}$ that correctly identifies pairs of images that show the same instance. We include the VQA prompt used during training in Fig. 18.
Reward Design. We define a simple binary outcome reward $R$ that compares the predicted final answer with the ground truth pairwise assignment: $R(y,\hat{y})=\mathbf{1}_{\{y=\hat{y}\}}$. Our supervision thus relies only on pairwise assignments and does not use any descriptive textual annotations.
Figure 9: Failure Cases on Fgvqa for models post-trained on Twin. For each example, we show the source dataset, images, and model predictions. Incorrect reasoning is highlighted in red. The red circle highlights hard-to-see objects.
Optimization. We optimize our VLM $\pi_{\theta}$ via GRPO [deepseekmath]. For each image pair $(I_{1},I_{2})$ with ground truth label $y$, we sample $G$ responses: $\mathcal{O}=\{(I_{1},I_{2},r^{(i)},\hat{y}^{(i)})\}^{G}_{i=1}$ and compute centralized advantages:
$A^{(i)}=R(y,\hat{y}^{(i)})-\frac{1}{G}\sum_{j=1}^{G}R(y,\hat{y}^{(j)})$ (1)
The GRPO objective then directly maximizes expected advantages while maintaining policy stability:
$\mathcal{L}(\theta)$ $=\frac{1}{G}\sum_{i=1}^{G}\Big(\hat{A}^{(i)}-\beta\,\text{KL}\big[\pi_{\theta}\,\big\|\,\pi_{\text{ref}}\big]\Big)$ (2)
$\hat{A}^{(i)}$ $=\min\Big(s^{(i)}A^{(i)},\,\text{clip}\big(s^{(i)},1-\epsilon,1+\epsilon\big)A^{(i)}\Big)$
where $s^{(i)}=\frac{\pi_{\theta}(r^{(i)},\hat{y}^{(i)}|I_{1},I_{2})}{\pi_{\text{old}}(r^{(i)},\hat{y}^{(i)}|I_{1},I_{2})}$ represents the importance ratio, $\epsilon=0.2$ is the clipping parameter, and $\beta=0.01$ is the KL penalty coefficient. The formulation encourages improved fine-grained understanding while simultaneously preventing drift from the pre-trained base policy $\pi_{\text{ref}}$.
Implementation. We train Qwen2.5-VL-3B-Instruct [qwen2.5vl] and InternVL3.5-1B-Instruct [internvl3] on Twin, chosen as leading open-source VLMs at the 3B and 1B scales. We do not freeze any part of the model, and train on $4$ A100 GPUs for 1 epoch. We use a batch size of $480$, group size $5$ and learning rate $10^{-6}$. We build on the verl [verl] repository for training. We detail all hyperparameters in Tab. 8.
C.2 Supervised Fine-Tuning
We compare post-training methods in Sec. 5.2. We provide additional details for supervised fine-tuning (SFT) used in those experiments. As our model produces both an explanation and final prediction, SFT requires supervision beyond the pairwise assignment labels used in RL. To obtain high-quality supervision, we prompt Gemini-2.5-Flash [gemini2.5] with samples from Twin and retain responses where Gemini is correct. We use default generation parameters for Gemini and collect a total of 136136K high-quality answers.
We train Qwen2.5-VL 3B end-to-end using SFT on the collected samples. We use LLaMa-Factory [llamafactory] for training, using a learning rate of $10^{-6}$ with a cosine scheduler. We train the SFT model for $2$ epochs, as this yielded better accuracy on Fgvqa. All hyperparameters are in Tab. 9.
Figure 10: Examples from Birds-to-Words [birdstowords]. The dataset features 3.33.3K pairs of different bird species, along with a human-annotated description of their differences. The differences are often coarse-grained (e.g. Pair 3, Pair 6 – which both display significant color differences), and the text descriptions subjective (e.g. Pair 4 – where animal 1 is described as ”more beautiful” than animal 2).
Appendix D Failure Cases
We include failure cases for models post-trained with Twin in Fig. 9. For each example, we show the source dataset, images, and model predictions. We find that although models post-trained with Twin demonstrate improved fine-grained understanding, they struggle with fine differences in color (e.g. in (a) where the ruins appears different colors due to lighting, and (f) where the watch accents differ between blue and gold). Moreover, extreme viewpoint variation, as seen in (b) remains a challenge. Lastly, we find that models struggle to extrapolate from incomplete views of the animals in (c) and (e). These challenging examples additionally highlight the difficulty of our new Fgvqa benchmark suite.
Figure 11: Examples from SpotTheDiff [spotthediff]. The dataset features 1313K pairs of images taken from different frames of the same video. Unlike Twin, that emphasizes fine differences in object texture and geometry, the differences in SpotTheDiff are primarily spatio-temporal (e.g. Pair 1 with the blue car leaving the frame, Pair 2 with the people moving) and highly coarse-grained (e.g. Pair 3).
Appendix E Comparisons to Prior Datasets
We provide a more detailed comparison to prior datasets Birds-To-Words [birdstowords] and SpotTheDiff [spotthediff]. While both of these datasets also feature pairs of images as input (similar to Twin), there are several key differences. First, the task in these datasets is to identify the differences between two images, with the assumption a priori that the images are indeed different. In contrast, Twin includes both negative pairs and positive pairs in which the instance is unchanged, requiring models to reason about both similarities and differences rather than only locating discrepancies. Second, the differences captured in Twin are substantially more fine-grained, often involving subtle variations in texture or part geometry that go beyond the relatively coarse differences in prior datasets. Third, our dataset is considerably larger in scale (561561K pairs vs. 1313K for SpotTheDiff and 3.33.3K for Birds-To-Words). We provide examples and describe the differences with each individual dataset in more detail below.
E.1 Birds-To-Words
Birds-to-words [birdstowords] features 3.33.3K pairs of images of different bird species, along with human-annotated descriptions of the differences. We show examples from the dataset in Fig. 10. We observe that, unlike the fine-grained differences in Twin, the differences are often much coarser (e.g. Pair 3 and Pair 6, which both feature significant color differences). Moreover, the human-annotated descriptions can be subjective (e.g. Pair 4 which describes one bird as “more beautiful”).
Compared to Twin, the dataset is orders of magnitude smaller (3.33.3K pairs vs the 561561K in Twin) and spans only bird species, while Twin features a wide range of object categories with significantly different geometries (e.g. earbuds and mugs). Additionally, Birds-to-Words features almost exclusively negative pairs. While a small number of pairs are annotated as depicting the same instance, these are anomalous cases and not a core aspect of the dataset – the task is fundamentally to describe the differences between two images. In contrast, Twin emphasizes both positive and negative pairs, requiring models to recognize not only differences but also similarities.
E.2 Comparison to SpotTheDiff
SpotTheDiff [spotthediff] comprises 1313K pairs of images sourced from two different frames of the same video. The videos used depict large scenes from security camera footage. As a result, the differences between frames are predominantly spatio-temporal, concerning which objects or people are visible or no longer visible in the scene (e.g. a car leaving the frame, a person moving positions). We visualize examples from SpotTheDiff [spotthediff] in Fig. 11. These examples highlight that differences are generally coarse-grained and, unlike the examples in Twin, do not require reasoning about viewpoint variation, part-level geometry, or fine-grained texture changes. Similar to Birds-to-Words, SpotTheDiff assumes the paired images are different, whereas Twin additionally emphasizes recognizing similarities through the inclusion of positive pairs. Finally, Twin is orders of magnitude larger in scale (561K vs. 13K pairs) and introduces challenges not present in SpotTheDiff, including substantial viewpoint variation, lighting changes, and differences in context and background.
Figure 12: Fgvqa Prompts: Twin-Eval.
Figure 13: Fgvqa Prompts: Ilias [ilias].
Figure 14: Fgvqa Prompts: Landmarks [landmarksv2].
Figure 15: Fgvqa Prompts: Met [met].
Figure 16: Fgvqa Prompts: Cub [cub].
Figure 17: Fgvqa Prompts: Inquire [inquire].
Figure 18: Prompt used for training with Twin.
Hyperparameter Value
Batch size 480
Group size 5
Max prompt length 2048
Max response length 2048
Learning rate 1e-6
Optimizer AdamW
Weight decay 0.01
KL coefficient 0.01
Clip ratio 0.2
Epochs 1
Rollout engine vLLM
Temperature 1.0
Top-p 1.0
Gradient clipping Max norm 1.0
Freeze vision tower False
Mixed precision bf16
Table 8: Hyperparameters for GRPO training.
Hyperparameter Value
Batch size 256
Max prompt length 2048
Max response length 2048
Learning rate 1e-6
Optimizer AdamW
Scheduler Cosine
Warmup ratio 0.1
Epochs 2
Deepspeed config ZeRO Stage 3
Gradient clipping auto
Freeze vision tower False
Mixed precision bf16
Table 9: Hyperparameters for SFT training.